# Privacy in Fine-tuning Large Language Models: Attacks, Defenses, and Future Directions

### v1

@@ 1 Introduction
Fine-tuning has emerged a highly popular approach for adapting pre-trained large language models (LLMs) to specific downstream tasks. By leveraging pre-trained knowledge and tailoring it to domain-specific needs, fine-tuning enables enhanced model performance on specialized applications. In recent years, numerous novel fine-tuning techniques have been developed [2, 15, 20, 22, 43]. Beyond the traditional full fine-tuning [6], methods like Parameter-Efficient Fine-Tuning (PEFT) [13] have gained traction, significantly reducing the resource overhead, such as time and computational cost. However, the rapid development of new fine-tuning methods and their diverse applications have also introduced new attack surfaces.
@@ 3 Privacy Attacks in Fine-tuning LLMs
  • Attacks aiming at manipulating model outputs: Often achieved through backdoor attacks, these involve adversaries embedding malicious triggers during the fine-tuning process to control the model’s behavior under specific conditions. Such attacks not only threaten the model’s security but also pose significant risks to its practical deployment.
@@ 5.3 Unique Characteristics of Fine-tuning Data
Fine-tuning datasets differ significantly from the general-purpose data used in pre-training, often exhibiting narrower domain coverage, task-specific distributions, and higher data sensitivity. These characteristics introduce unique privacy risks. Narrower and task-specific distributions make fine-tuning datasets more vulnerable to targeted attacks. Smaller dataset sizes exacerbate memorization issues, increasing the risk of sensitive information being leaked. However, there is currently a lack of privacy research specifically focused on the characteristics of fine-tuning data. To address this issue, future research can be conducted from both the attack and defense perspectives. From the perspective of attack, exploring domain-specific or task-specific attacks can help us understand the unique traits of fine-tuning datasets and design more effective defenses. From the perspective of defense, we need to explore task-specific or domain-specific privacy mechanisms to adapt to more specific tasks.
@@ 5.5 Balancing Utility and Privacy in Fine-tuning Defenses
The utility-privacy trade-off remains a longstanding challenge in fine-tuning large language models (LLMs). In fact, most defense methods also consider their utility-privacy trade-off as a critical factor. This trade-off, a well-known challenge in privacy protection, continues to be relevant in the fine-tuning stage. Moreover, given the high performance demands of fine-tuning for downstream tasks, maintaining utility becomes even more critical and cannot be overlooked. To address that, dynamic adjustment defense mechanisms or task-aware defense strategies can be potential directions.

### v2

@@ 1 Introduction
Fine-tuning has emerged a highly popular approach for adapting pre-trained large language models (LLMs) to specific downstream tasks. By leveraging pre-trained knowledge and tailoring it to domain-specific needs, fine-tuning enables enhanced model performance on specialized applications and domains like autonomous systems, legal applications, and healthcare. However, this process often involves sensitive information, such as sensor data, confidential legal documents, or personal medical records, raising significant privacy concerns. In recent years, numerous novel fine-tuning techniques have been developed [2, 15, 20, 22, 43]. Beyond the traditional full fine-tuning [6], methods like Parameter-Efficient Fine-Tuning (PEFT) [13] have gained traction, significantly reducing the resource overhead, such as time and computational cost. Conversely, the rapid development of new fine-tuning methods and their diverse applications have also introduced new attack surfaces.
@@ 3 Privacy Attacks in Fine-tuning LLMs
  • Attacks aiming at manipulating model outputs: Often achieved through backdoor attacks, these involve adversaries embedding malicious triggers during the fine-tuning process to control the model’s behavior under specific conditions. These attacks can be exploited to induce the model to output sensitive information, and they can be combined with other attack methods to further enhance the overall success rate.
@@
Furthermore, we intend to provide a more detailed explanation of the content within the Capability. Reference Model means the attacker needs the original model or a fine-tuned one as a reference in the attack. Fine-tuning Access indicates that an attacker needs to have the access to fine-tune the model. Poison Injection while Pre-training indicates that an attacker has the capability to inject malicious data during the model’s pre-training phase. Publishing Models refers to the capability of publicly releasing and sharing models. Masked Fine-tuning Data means an attacker can get fine-tuning datasets with the sensitive information masked.
@@ 5.3 Unique Characteristics of Fine-tuning Data
Fine-tuning datasets differ significantly from the general-purpose data used in pre-training, often exhibiting narrower domain coverage, task-specific distributions, and higher data sensitivity. For instance, a fine-tuning dataset built from medical records or legal documents typically contains specialized terminology and sensitive personal information, unlike the vast and heterogeneous corpora used in pre-training. These characteristics introduce unique privacy risks. Narrower and task-specific distributions make fine-tuning datasets more vulnerable to targeted attacks. A fine-tuning dataset for legal document summarization might inadvertently retain verbatim excerpts from confidential case files, which could then be extracted by extraction attacks. However, there is currently a lack of privacy research specifically focused on the characteristics of fine-tuning data. To address this issue, future research can be conducted from both the attack and defense perspectives. From the perspective of attack, exploring domain-specific or task-specific attacks can help us understand the unique traits of fine-tuning datasets and design more effective defenses. From the perspective of defense, we need to explore task-specific or domain-specific privacy mechanisms to adapt to more specific tasks.
@@ 5.5 Balancing Utility and Privacy in Fine-tuning Defenses
The utility-privacy trade-off remains a longstanding challenge in fine-tuning large language models (LLMs). While defense methods aim to mitigate privacy risks, preserving model performance for downstream tasks remains equally critical. In fact, most defense methods also consider their utility-privacy trade-off as a critical factor. This trade-off, a well-known challenge in privacy protection, continues to be relevant in the fine-tuning stage. Moreover, given the high performance demands of fine-tuning for downstream tasks, maintaining utility becomes even more critical and cannot be overlooked. To address that, dynamic adjustment defense mechanisms or task-aware defense strategies can be potential directions.
@@ 6 Conclusion
7 Acknowledgment
This work is partially support by JSPS KAKENHI JP23K24851, JST PRESTO JPMJPR23P5, JST CREST JPMJCR21M2, JSPS KAKENHI Grant Number JP24H00685. Lei Chen’s work is partially supported by National Key Research and Development Program of China Grant No. 2023YFF0725100, National Science Foundation of China (NSFC) under Grant No. U22B2060, Guangdong-Hong Kong Technology Innovation Joint Funding Scheme Project No. 2024-A0505040012, the Hong Kong RGC GRF Project 16213620, RIF Project R6020-19, AOE Project AoE/E-603/18, Theme-based project TRS T41-603/20R, CRF Project C2004-21G, Guangdong Province Science and Technology Plan Project 2023A0505030011, Guangzhou municipality big data intelligence key lab, 2023-A03J0012, Hong Kong ITC ITF grants MHX/078/21 and PRP/004/22FX, Zhujiang scholar program 2021JC02X170, Microsoft Research Asia Collaborative Research Grant, HKUST-Webank joint research lab and 2023 HKUST Shenzhen-Hong Kong Collaborative Innovation Institute Green Sustainability Special Fund, from Shui On Xintiandi and the InnoSpace GBA.